OpenStudy (anonymous):

how do you convert y=2x-1 to polar form???

OpenStudy (jdoe0001):

\(\bf \textit{keep in mind that }x=rcos(\theta)\qquad y=rsin(\theta)\qquad thus\\ \quad \\ y=2x-1\implies rsin(\theta)=2[rcos(\theta)]-1\) solve for "r"

OpenStudy (anonymous):

so i finished the problem but I'm not sure if it is right. is the answer 1/2cos-sin=r???

OpenStudy (jdoe0001):

1/2cos-sin=r \(\checkmark\)

OpenStudy (jdoe0001):

yes \(\bf \cfrac{1}{2cos(\theta)-sin(\theta)}=r\)

OpenStudy (anonymous):

thanks a lot

OpenStudy (jdoe0001):

yw